Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

Sales $ 20,000
Variable expenses 13,000
Contribution margin 7,000
Fixed expenses 3,780
Net operating income $ 3,220

7. If the variable cost per unit increases by $1, spending on advertising increases by $1,100, and unit sales increase by 120 units, what would be the net operating income?

8. What is the break-even point in unit sales

13. Using the degree of operating leverage, what is the estimated percent increase in net operating income of a 5% increase in sales? (Round your intermediate calculations and final answer to 2 decimal places.)

14. Assume that the amounts of the company’s total variable expenses and total fixed expenses were reversed. In other words, assume that the total variable expenses are $3,780 and the total fixed expenses are $13,000. Under this scenario and assuming that total sales remain the same, what is the degree of operating leverage? (Round your answer to 2 decimal places.)

15. Assume that the amounts of the company’s total variable expenses and total fixed expenses were reversed. In other words, assume that the total variable expenses are $3,780 and the total fixed expenses are $13,000. Given this scenario and assuming that total sales remain the same. Using the degree of calculated operating leverage, what is the estimated percent increase in net operating income of a 5% increase in sales? (Round your intermediate calculations and final answer to 2 decimal places.)
Answer #1

7) Calculate net operating income

Sales (1120*20) 22400
Variable cost (1120*14) 15680
Contribution margin 6720
Fixed cost (3780+1100) 4880
Net income 1840

8) Break even unit = 3780/7 = 540 Units

13) Degree of operating leverage = 7000/3220 = 2.17

Percentage increase in net income = 2.17*5 = 10.85%

14) Contribution margin (in different scenario)

Sales 20000
Variable cost 3780
Contribution margin 16220
Fixed cost 13000
Net income 3220

Degree of operating leverage = 16220/3220 = 5.04

15) Increase in net income = 5.04*5 = 25.20%